OpenStudy (anonymous):

What's the minimum or maximum value of the function f(x) = x2 - 6x + 3.

OpenStudy (anonymous):

minimum is the second coordinate of the vertex. first coordinate is always \(-\frac{b}{2a}\) in your example \(-\frac{-6}{2\times 1}=3\) replace \(x\) by 3 and see what you get

OpenStudy (anonymous):

Would it be -6?

OpenStudy (anonymous):

bingo!

OpenStudy (anonymous):

Thank you